### v2

# Establishing Linear Surrogate Regret Bounds for Convex Smooth Losses via Convolutional Fenchel–Young Losses

Abstract
Surrogate regret bounds, also known as excess risk bounds, bridge the gap between the convergence rates of surrogate and target losses, with linear bounds favorable for their lossless regret transfer. While convex smooth surrogate losses are appealing in particular due to the efficient estimation and optimization, the existence of a trade-off between the smoothness and linear regret bound has been believed in the community. That being said, the better optimization and estimation properties of convex smooth surrogate losses may inevitably deteriorate after undergoing the regret transfer onto a target loss. We overcome this dilemma for arbitrary discrete target losses by constructing a convex smooth surrogate loss, which entails a linear surrogate regret bound composed with a tailored prediction link. The construction is based on Fenchel–Young losses generated by the convolutional negentropy, which are equivalent to the infimal convolution of a generalized negentropy and the target Bayes risk. Consequently, the infimal convolution enables us to derive a smooth loss while maintaining the surrogate regret bound linear. We additionally benefit from the infimal convolution to have a consistent estimator of the underlying class probability. Our results are overall a novel demonstration of how convex analysis penetrates into optimization and statistical efficiency in risk minimization.
1 Introduction
The risk of a machine learning model is often measured by the expectation of a target loss $\ell$ that quantifies the error between the model prediction $t$ and a natural label $y$. However, minimizing such a target risk over a dataset is often computationally hard because a target prediction problem is usually discretely structured, including multiclass, multilabel, top-$k$ prediction problems. For this reason, a tractable surrogate risk induced by a surrogate loss $L(\bm{\theta},y)$ serves as an essential proxy, where $\bm{\theta}$ is a score vector that often takes value in $\mathbb{R}^{\varrho}$.
An ideal surrogate loss should be convex, smooth, and calibrated toward a given target prediction problem. Convexity and smoothness (or Lipschitz continuous gradient) of a loss function have been fundamental properties both in optimization and statistical estimation—indeed, classical optimization theory reveals that convex and smooth functions can be efficiently optimized with first-order methods [47], which is prevalent in modern machine learning. In addition, several studies have demonstrated that the convexity and smoothness can further enhance fast rates in the risk estimation [63, 71]. Meanwhile, calibration is regarded as a minimal requirement on surrogate losses, ensuring that the surrogate risk minimization leads to the target risk minimization [6, 64].111 Readers should distinguish the calibration property in this context from the calibration of probabilistic predictions [24]. To establish a relationship between surrogate and target risks, surrogate regret bounds play a crucial role. Therein the regret (or the suboptimality) of a target loss is controlled by that of a surrogate loss through a non-decreasing rate function $\psi:\mathbb{R}_{\geq 0}\to\mathbb{R}_{\geq 0}$. A surrogate regret bound can be informally written as follows: for any score vector $\bm{\theta}\in\mathbb{R}^{\varrho}$ and a class distribution $\bm{\eta}$,
$\mathtt{Regret}_{\ell}(\varphi(\bm{\theta}),\bm{\eta})\leq\psi\left(\mathtt{Regret}_{L}(\bm{\theta},\bm{\eta})\right),$ (1)
where $\varphi$ is a prediction link that converts a score vector $\bm{\theta}$ into a target prediction $\varphi(\bm{\theta})$. If the regret rate function is linear $\psi(r)=\mathcal{O}(r)$, which is the best possible (data-independent) regret rate,222 A super-linear regret bound is possible if we allow a surrogate loss to depend on a data distribution [70], which is not considered here. then the optimization and estimation errors of the surrogate regret are optimally translated to the target regret. For example, under the binary classification target loss, Bartlett et al. [6] demonstrated that the linear regret convergence is possible when the hinge loss is used as a surrogate loss; additionally, symmetric losses [15] and polyhedral losses [25] yield the linear regret convergence. Unfortunately, these loss functions lack either convexity or smoothness, which deteriorates the optimization and estimation errors of the surrogate regret, even if they enjoy linear regret bounds. By contrast, a square-root regret rate function $\psi(r)=\mathcal{O}(\sqrt{r})$ is commonly encountered when it comes to convex smooth surrogate losses, such as the logistic, exponential, and squared losses [6, 49, 25]. These loss functions enjoy better optimization and estimation properties, yet suffer from larger target regrets due to the suboptimal regret rate. Therefore, it has been open for a long time to develop a convex smooth surrogate loss without sacrificing the regret convergence rate.
Notwithstanding, the previous literature in this line has implied that such an ideal surrogate loss may be inconceivable. Mahdavi et al. [33] considered this question with an interpolated loss between the hinge (non-smooth) and logistic (smooth) losses and showed that we inevitably face the trade-off between generalization and optimization unless strong distributional assumptions are imposed. Further, Frongillo and Waggoner [25] proved that locally smooth and strongly convex losses must suffer from a square-root regret convergence rate at least. Ramaswamy et al. [53] blame convex smooth losses for redundantly modeling continuous class distributions, which is unnecessary if our goal is merely to solve discrete target problems.
In this paper, we demonstrate that this seemingly impossible trade-off can be overcome for every discrete target loss $\ell$. Specifically, we build a convex smooth surrogate loss built upon the framework of Fenchel–Young losses [10]—a framework to generate a loss function from a generalized negentropy and its conjugate. In a nutshell, our main results are summarized as follows:
Theorem 1 (Informal version of Theorem 15)
For any discrete target loss $\ell$, there exist a convex smooth surrogate loss $L$ (defined over a score $\bm{\theta}\in\mathbb{R}^{\varrho}$) and prediction link $\varphi$ such that the surrogate regret bound (1) holds with some linear rate $\psi(r)=\mathcal{O}(r)$.
This existence is proved constructively, which provides a systematic framework on the construction of convex smooth surrogate losses and their corresponding prediction links with linear surrogate regret bounds. Our high-level construction, which significantly leverages convex analysis and is detailed in Section 3.1, proceeds as follows. First, a user chooses a strongly convex negentropy with some regularity conditions, such as the Shannon negentropy. We encode the structure of a target loss $\ell$ into the chosen base negentropy by adding the negative Bayes risk of $\ell$. This additivity eventually translates into the infimal convolution in the induced Fenchel–Young loss, which we call the convolutional Fenchel–Young loss. The convolutional Fenchel–Young loss is endowed with the convexity and smoothness arising from the base negentropy, shown in Section 3.2. Then we can obtain a prediction link via the infimal convolution. Paired with the convolutional Fenchel–Young loss, it admits a linear surrogate regret bound, demonstrated in Section 3.3.
In addition, we improve the multiplicative term in the initial linear surrogate regret bound in Section 3.4 to make the bound tighter, which exploits the low-rank structure of a target loss $\ell$. As a by-product of the loss smoothness, we can provide a Fisher-consistent probability estimator of the underlying probability in Section 3.5. Finally, Section 4 instantiates the framework of the convolutional Fenchel–Young loss for the multiclass classification problem, highlighting the efficient computation of the prediction link.
1.1 Related Works
Surrogate losses for general discrete prediction problems.
A discrete prediction problem aims to predict $t$ that minimizes a discrete target loss $\ell(t,y)$ over the class distribution, and the study of convex surrogates for $\ell$ is of significant interest. Extensive research has been conducted on surrogate losses for specific discrete tasks, including but not limited to classification [6, 74, 65, 67, 61, 46, 43], top-$k$ [69, 66], and multilabel learning [26, 73, 37, 32, 44].
In contrast to ad-hoc approaches, recent studies have advanced the principled design of calibrated convex surrogate losses for general discrete prediction problems, without imposing restrictions on the target losses. In Ramaswamy and Agarwal [51] and Ramaswamy et al. [54], surrogate losses based on the squared loss and error correcting output codes are proposed, respectively, which embed the structure of a target problem into a surrogate loss and are shown to be calibrated with the corresponding target losses. The design of calibrated surrogate losses has also been extensively studied in the context of structured prediction [19, 52, 17, 18, 50, 48, 49], where the structural/low-rank properties of the target losses are exploited to construct more efficient surrogates. As opposed to the smooth losses utilized in the works above, polyhedral losses [21, 22, 23] has attracted attention recently, which provides a systematic framework for constructing efficiently calibrated yet non-smooth convex losses based on a discrete target loss.
Previous studies on surrogate regret bounds.
Surrogate regret bounds have been well studied for margin losses under binary and multiclass classification [74, 6, 65, 60, 39, 62], where the target loss is the 0-1 loss. Similarly, proper losses [14, 27, 56, 2, 29, 45] have been shown to provide surrogate regret bounds w.r.t. the $L_{1}$ distance between the estimated and true class probabilities [55, 4, 5], which facilitates analyses for downstream tasks. For structured prediction, surrogate regret bounds are sometimes called comparison inequalities [17, 50, 48, 8], which are typically of square-root type. Notably, Frongillo and Waggoner [25] demonstrated that polyhedral losses exhibit linear surrogate regret bounds for general discrete prediction problems, covering various piecewise linear non-smooth losses [69, 53] as special cases. However, they lack the smoothness. To the contrary, Mao et al. [36] and Mao et al. [35] propose smooth losses with linear regret rates but lacking convexity.
Whereas a growing line of research has focused on $\mathcal{H}$-consistency to analyze how the restriction to a function class $\mathcal{H}$ affects consistency [31, 72, 3], we adopt surrogate regrets over all measurable score functions because the analysis is often more transparent and it is reasonable to suppose that the function class is sufficiently expressive under the overparametrization regime. The extension to $\mathcal{H}$-consistency is rather straightforward by integrating the minimizability gap [35, 38, 37, 40, 34].
2 Preliminaries
Let $[d]\coloneqq\{1,2,\cdots,d\}$ and [[]]delimited-[]delimited-[][\![$\cdot$]\!][ [ ⋅ ] ] is the Iverson bracket. The $p$-norm is denoted by $\|\cdot\|_{p}$, which we assume to be the $2$-norm unless otherwise noted. Let $\overline{\mathbb{R}}\coloneqq\mathbb{R}\cup\{\infty\}$ be the extended real-line and $\Delta^{d}\coloneqq\{\bm{\eta}\in\mathbb{R}^{d}_{\geq 0}:\|\bm{\eta}\|_{1}=1\}$ the $d$-simplex. For a set $\mathcal{S}\subseteq\mathbb{R}^{d}$, $\mathrm{int}(\mathcal{S})$ and $\mathrm{relint}(\mathcal{S})$ are its interior and relative interior, respectively, and $\mathrm{conv}(S)$ is its convex hull. The indicator function is denoted by $\mathbb{I}_{\mathcal{S}}:\mathbb{R}^{d}\to\{0,+\infty\}$, where $\mathbb{I}_{\mathcal{S}}(\bm{\theta})=0$ if $\bm{\theta}\in\mathcal{S}$ and +$\infty$+ ∞ otherwise. For a function $f:\mathbb{R}^{d}\to\overline{\mathbb{R}}$, $\mathrm{dom}(\Omega)\coloneqq\{\bm{\theta}\in\mathbb{R}^{d}:\Omega(\bm{\theta})<+\infty\}$ is its effective domain. A function $f$ is extended to be set-valued with slight abuse of notation by $f(\mathcal{S})\coloneqq\{f(s):s\in\mathcal{S}\}$. The Fenchel conjugate of $\Omega$ is $\Omega^{*}(\bm{\theta})\coloneqq\sup_{\bm{p}\in\mathrm{dom}(\Omega)}\bm{\theta}^{\top}\bm{p}-\Omega(\bm{p})$. The canonical basis of the Euclidean space is denoted by $\{\bm{e}_{i}\}$, where the dimensionality depends on the contexts.
2.1 Discrete Prediction Problems and Target Losses
In this section, we review the target losses that define discrete prediction problems, and further introduce their equivalent lower-dimensional decompositions.
Discrete prediction problem:
Let $\mathcal{Y}=[K]$ be the finite class space. The class distribution on $\mathcal{Y}$ is $\bm{\eta}\in\Delta^{K}$ such that class $Y=y$ with probability $\eta_{y}$. A discrete prediction problem aims to find a target prediction $t$ from the finite prediction space $\widehat{\mathcal{Y}}\coloneqq[N]$ for each $\bm{\eta}$ that minimizes the expectation of a discrete target loss $\ell:\widehat{\mathcal{Y}}\times\mathcal{Y}\to\mathbb{R}$ over $y\sim\bm{\eta}$. This expectation is called the risk w.r.t. the target loss $\ell$: $R_{\ell}(t,\bm{\eta})\coloneqq\mathbb{E}_{y\sim\bm{\eta}}[\ell(t,y)]=\langle\bm{\eta},\bm{\ell}(t)\rangle$, where $\bm{\ell}(t)\in\mathbb{R}^{K}$ is the loss vector such that $\bm{\ell}(t)_{y}=\ell(t,y)$. The Bayes risk of $\ell$ at $\bm{\eta}$ is $\underline{R}_{\ell}(\bm{\eta})\coloneqq\min_{t\in\widehat{\mathcal{Y}}}R_{\ell}(t,\bm{\eta})$, which is a concave function of $\bm{\eta}$. The suboptimality of a prediction $t$ w.r.t. the target loss $\ell$ over a class distribution $\bm{\eta}$ is characterized by the target regret, which is the gap between its risk and the Bayes risk.
Definition 2 (Target regret)
Given a discrete target loss $\ell$, the target regret of a prediction $t$ w.r.t. a class probability $\bm{\eta}\in\Delta^{K}$ is defined as follows:
$\mathtt{Regret}_{\ell}(t,\bm{\eta})\coloneqq R_{\ell}(t,\bm{\eta})-\underline{R}_{\ell}(\bm{\eta}).$ (2)
Equivalent lower-dimensional decomposition.
By encoding every possible class label $y\in\mathcal{Y}$ into the canonical basis $\bm{e}_{y}\in\mathbb{R}^{K}$, we can express any target loss in the following form:
$\ell(t,y)=\langle\bm{e}_{y},\bm{\ell}(t)\rangle.$ (3)
An equivalent but more efficient decomposition of (3) called Structure Encoding Loss Functions (SELF) is introduced in Ciliberto et al. [17], which allows lower-dimensional encoding of class labels. It is often leveraged by follow-up works to induce efficient loss functions, and is further generalized by the Affine Decomposition in Blondel [8, (12)]. In this work, we will also represent discrete target losses in this decomposed form.
Definition 3 ($(\bm{\rho},\bm{\ell}^{\bm{\rho}})$-decomposition)
For a discrete target loss $\ell:\widehat{\mathcal{Y}}\times\mathcal{Y}\to\mathbb{R}$, its $(\bm{\rho},\bm{\ell}^{\bm{\rho}})$-decomposition is given as follows:
$\ell(t,y)=\langle\bm{\rho}(y),\bm{\ell}^{\bm{\rho}}(t)\rangle+c(y),$ (4)
where $\bm{\rho}:\mathcal{Y}\to\mathbb{R}^{\varrho}$ is a label encoding function that maps discrete labels into the $\varrho$-dimensional Euclidean space, $\bm{\ell}^{\bm{\rho}}:\widehat{\mathcal{Y}}\to\mathbb{R}^{\varrho}$ is the corresponding  loss encoding function, and the term $c:\mathcal{Y}\to\mathbb{R}$ is the remainder independent of prediction $t$.
This loss decomposition contains the Affine Decomposition [8, (12)], which generalizes SELF [17]. Any discrete target loss enjoys this decomposition since it can be trivially decomposed by $\bm{\rho}(y)=\bm{e}_{y}$, $\bm{\ell}^{\bm{\rho}}(t)=\bm{\ell}(t)$, and $c=0$ with $\varrho=K$, which immediately recovers (3). With a properly chosen $\bm{\rho}$, the encoded cardinality $\varrho$ can often be greatly smaller than $K$, particularly in the context of structured prediction. For example, consider multilabel classification, where $\mathcal{Y}=\widehat{\mathcal{Y}}=[K]$, $K=2^{\varrho}$, and each $y\in[K]$ corresponds to a unique binary vector $\bm{\nu}(y)\in\{0,1\}^{\varrho}$ representing a set of binary labels. Suppose our target loss is the Hamming loss $\ell_{H}(t,y)\coloneqq\bm{1}^{\top}(\bm{\nu}(t)+\bm{\nu}(y))-2\langle\bm{\nu}(t),\bm{\nu}(y)\rangle$, which is the Hamming distance between $\bm{\nu}(t)$ and $\bm{\nu}(y)$. This entails the decomposition $\bm{\rho}(y)=\bm{1}-2\bm{\nu}(y)$, $\bm{\ell}^{\bm{\rho}}(t)=\bm{\nu}(t)$, and $c(y)=\bm{1}^{\top}\bm{\nu}(y)$. Here, $\varrho=\ln_{2}K$ significantly reduces the dimensionality from the cardinality of $\mathcal{Y}$. We refer reader to Blondel [8, Appendix A] for more examples of discrete prediction problems. The cardinality $\varrho$ of a label encoding function $\bm{\rho}$ is closely related to surrogate regret bounds later in Section 3.4.
2.2 Surrogate Losses and Surrogate Regret Bounds
Unlike discrete target losses assessing a discrete prediction $t\in\widehat{\mathcal{Y}}$, a surrogate loss $L:\mathbb{R}^{d}\times\mathcal{Y}\to\overline{\mathbb{R}}$ receives a continuous score $\bm{\theta}\in\mathbb{R}^{d}$. Then a prediction link $\varphi:\mathbb{R}^{d}\to\widehat{\mathcal{Y}}$ is used to transform a score $\bm{\theta}$ to a prediction $t$. Its risk and Bayes risk are defined as $R_{L}(\bm{\theta},\bm{\eta})\coloneqq\mathbb{E}_{y\sim\bm{\eta}}[L(\bm{\theta},y)]=\langle\bm{\eta},\bm{L}(\bm{\theta})\rangle$ and $\underline{R}_{L}(\bm{\eta})\coloneqq\inf_{\bm{\theta}\in\mathbb{R}^{d}}R_{L}(\bm{\theta},\bm{\eta})$, respectively, where $\bm{L}(\bm{\theta})\coloneqq[L(\bm{\theta},y)]_{y=1}^{K}$. The regret of a surrogate loss is defined as the gap between the risk of a score $\bm{\theta}$ and the Bayes risk, similar to the target regret.
Definition 4 (Surrogate regret)
Given a surrogate loss $L$, the surrogate regret of score $\bm{\theta}$ w.r.t. a class probability $\bm{\eta}\in\Delta^{K}$ is defined as follows:
$\mathtt{Regret}_{L}(\bm{\theta},\bm{\eta})\coloneqq R_{L}(\bm{\theta},\bm{\eta})-\underline{R}_{L}(\bm{\eta}).$ (5)
For a desirable surrogate loss, the convergence of its surrogate regret will dominate the target regret. That is, for a fixed class distribution $\bm{\eta}$, a convergent $\bm{\theta}$ such that $\mathtt{Regret}_{L}(\bm{\theta},\bm{\eta})\to 0$ should imply $\mathtt{Regret}_{\ell}(\varphi(\bm{\theta}),\bm{\eta})\to 0$. This convergence relationship indicates that the target regret minimization can be achieved by the minimization of the surrogate regret adopted with an appropriate prediction link. A surrogate regret bound offers a quantitative characterization of this relationship through a regret rate function $\psi$, which is the key focus of this work.
Definition 5 (Surrogate regret bound)
A surrogate loss $L$ and prediction link $\varphi$ entail a surrogate regret bound w.r.t. target $\ell$ with a non-decreasing regret rate function $\psi:\mathbb{R}_{\geq 0}\to\mathbb{R}_{\geq 0}$ satisfying $\psi(0)=0$ if the following inequality holds:
$\mathtt{Regret}_{\ell}(\varphi(\bm{\theta}),\bm{\eta})\leq\psi(\mathtt{Regret}_{L}(\bm{\theta},\bm{\eta})),~{}~{}~{}~{}~{}\text{for any~{}}(\bm{\theta},\bm{\eta})\in\mathbb{R}^{\varrho}\times\Delta^{K}.$ (6)
While a linear regret rate $\psi(r)=\mathcal{O}(r)$ is the best possible, previously known examples are either non-smooth or non-convex surrogate losses, including the (non-smooth) hinge loss [6] and (non-convex) sigmoid loss [15]. These loss functions may face challenges in optimization and estimation [63]. In this work, we aim to develop a framework that facilitates the use of convex smooth surrogates with linear surrogate regret bounds.
2.3 Fenchel–Young Loss
In this work, we build upon Fenchel–Young losses [10] to construct convex smooth surrogates equipped with linear regret rate functions. Let us review Fenchel–Young losses first.
Definition 6 (Fenchel–Young loss)
For $\Omega:\mathbb{R}^{d}\to\overline{\mathbb{R}}$, the associated Fenchel–Young loss $L_{\Omega}:\mathrm{dom}(\Omega^{*})\times\mathrm{dom}(\Omega)\to\mathbb{R}_{\geq 0}$ is defined as follows:
$L_{\Omega}(\bm{\theta},\bm{p})=\Omega(\bm{p})+\Omega^{*}(\bm{\theta})-\langle\bm{\theta},\bm{p}\rangle.$ (7)
The function $\Omega$ is often regarded as a negentropy, which equals the Bayes risk of its induced Fenchel–Young loss. Fenchel–Young losses often impose additional requirements on the domains of $\Omega$ and $\Omega^{*}$, tailored to specific target problems, and we will highlight them when necessary.
Fenchel–Young losses possess favorable properties: they are inherently convex (in score $\bm{\theta}$) by definition. Moreover, they can induce Fisher-consistent estimators of $\mathbb{E}_{Y\sim\bm{\eta}}[\bm{\rho}(Y)]$, which recovers the class distribution $\bm{\eta}$ under multiclass classification with $\bm{\rho}(y)=\bm{e}_{y}$; or the linear property in general [1]. Many common losses are encompassed in Fenchel–Young losses, including the cross-entropy loss, squared loss, and Crammer–Singer loss [20]. For example, $\Omega(\bm{p})=\frac{1}{2}\|\bm{p}\|^{2}+\mathbb{I}_{\Delta^{K}}(\bm{p})$ generates the sparsemax loss, which induces the sparsemax as a Fisher-consistent probability estimator of $\bm{\eta}$ [41].
3 Convex Smooth Surrogates with Linear Surrogate Regret Bounds
In this section, we first show our construction of surrogate loss and prediction links. After showing that this loss is convex and smooth with an appropriately chosen base negentropy (Theorem 11), we move on to the main result of this paper, linear surrogate regret bounds with convex smooth surrogate losses (Theorem 13). We further discuss the computational aspects and improve the regret bound constant. The missing proofs of technical results are deferred to Appendix A.
3.1 Construction
We design a convex smooth surrogate loss built upon the framework of Fenchel–Young losses. To make its surrogate regret bound linear, we craft an negentropy function by delicately leveraging the structure of a target prediction problem. Denote by $T$ the polyhedral convex function $T(\bm{p})=-\min_{t\in\widehat{\mathcal{Y}}}\left\langle\bm{p},\bm{\ell}^{\bm{\rho}}(t)\right\rangle$. Then, we have the following definition:
Definition 7 (Convolutional negentropy)
Suppose $(\bm{\rho},\bm{\ell}^{\bm{\rho}})$-decomposition for a target loss $\ell$. For an negentropy function $\Omega:\mathbb{R}^{\varrho}\to\overline{\mathbb{R}}$, its convolutional negentropy is defined as follows:
$\Omega_{T}(\bm{p})=\Omega(\bm{p})+T(\bm{p})$ (8)
While commonly used negentropy functions in Fenchel–Young losses, such as the Shannon negentropy and norm negentropy [12], are unstructured toward a target loss, the convolutional negentropy $\Omega_{T}$ encodes a target loss $\ell$ explicitly into the base negentropy $\Omega$ via $T$. This polyhedral convex function $T$ is an affinely transformed Bayes risk of a target loss $\ell$. Indeed, with the linear property $\bm{p}=\mathbb{E}_{y\sim\bm{\eta}}[\bm{\rho}(y)]$, we have
$T(\bm{p})=-\min_{t\in\widehat{\mathcal{Y}}}\langle\bm{p},\bm{\ell}^{\bm{\rho}}(t)\rangle=-\min_{t\in\widehat{\mathcal{Y}}}\mathbb{E}_{y\sim\bm{\eta}}[\langle\bm{\rho}(y),\bm{\ell}^{\bm{\rho}}(t)\rangle]=\sum_{y\in[K]}\eta_{y}c(y)-\underline{R}_{\ell}(\bm{\eta}).$ (9)
The base negentropy $\Omega$ is up to our choice. Before deriving the conjugate of $\Omega_{T}$, we make the following requirement on $\Omega$ throughout this work for a well-behaved convolutional negentropy.
Condition 1
$\Omega$ is proper convex and lower-semicontinuous (l.s.c.), and satisfies $\mathrm{conv}(\bm{\rho}(\mathcal{Y}))\subseteq\mathrm{dom}(\Omega)$ and $\mathrm{dom}(\Omega^{*})=\mathbb{R}^{\varrho}$.
It is a mild requirement. Indeed, proper convexity and lower-semicontinuity merely aim to avoid pathologies, and the domain assumptions are met by a differentiable and finite $\Omega$ over the valid prediction space $\mathrm{conv}(\bm{\rho}(\mathcal{Y}))$. The domain assumption on $\Omega^{*}$ is crucial for the induced loss to have an unconstrained domain $\mathbb{R}^{\varrho}$. Then we show an explicit form of the conjugated convolutional negentropy.
Lemma 8 (Conjugate of $\Omega_{T}$)
Suppose Condition 1 holds, then $\Omega_{T}^{*}$ is proper convex and l.s.c. with $\mathrm{dom}(\Omega_{T}^{*})=\mathbb{R}^{\varrho}$, and can be expressed as follows:
$\Omega_{T}^{*}(\bm{\theta})=\inf_{\bm{\pi}\in\Delta^{N}}\Omega^{*}(\bm{\theta}+\mathcal{L}^{\bm{\rho}}\bm{\pi})\quad\text{for any $\bm{\theta}\in\mathbb{R}^{\varrho}$},$ (10)
where $\mathcal{L}^{\bm{\rho}}\in\mathbb{R}^{\varrho\times N}$ is the loss matrix with the $t$-th column being $\bm{\ell}^{\bm{\rho}}(t)\in\mathbb{R}^{\varrho}$ for $t\in\widehat{\mathcal{Y}}$. In addition, there always exists $\bm{\pi}\in\Delta^{N}$ that achieves the infimum of (10) for any $\bm{\theta}\in\mathbb{R}^{\varrho}$.
The conjugated form (10) owes to the infimal convolution between the base negentropy $\Omega$ and the transformed target Bayes risk $\min_{t\in\widehat{\mathcal{Y}}}\langle\bm{p},\bm{\ell}^{\bm{\rho}}(t)\rangle$. We coined the name of the convolutional negentropy inspired by this structure. Intuitively, the conjugated convolutional negentropy $\Omega_{T}^{*}$ can be viewed as a “perturbed” version of the conjugated base negentropy $\Omega^{*}$ toward the direction of the target loss matrix $\mathcal{L}^{\bm{\rho}}$. This result facilitates a more concrete formulation of the Fenchel–Young loss generated by $\Omega_{T}$, which we refer to as the convolutional Fenchel–Young loss—the central focus of this work.
Definition 9 (Convolutional Fenchel–Young loss)
Suppose that a discrete target loss $\ell$ enjoys $(\bm{\rho},\bm{\ell}^{\bm{\rho}})$-decomposition. For an negentropy function $\Omega:\mathbb{R}^{\varrho}\to\overline{\mathbb{R}}$ satisfying Condition 1, the surrogate loss $L_{\Omega_{T}}:\mathbb{R}^{\varrho}\times\mathcal{Y}\to\overline{\mathbb{R}}$ induced by the convolutional negentropy $\Omega_{T}$ is defined as follows:
$L_{\Omega_{T}}(\bm{\theta},y)=\min_{\bm{\pi}\in\Delta^{N}}\Omega^{*}(\bm{\theta}+\mathcal{L}^{\bm{\rho}}\bm{\pi})+\Omega_{T}(\bm{\rho}(y))-\langle\bm{\theta},\bm{\rho}(y)\rangle.$ (11)
Note that $L_{\Omega_{T}}(\bm{\theta},\cdot)$ is deliberately constrained to the class space $\mathcal{Y}$ instead of the general domain $\mathrm{dom}(\Omega_{T})$, to align with the surrogate loss form introduced in Section 2.2.
Given a surrogate loss, we need to specify a prediction link. For general discrete prediction problems where $\mathcal{Y}\neq\widehat{\mathcal{Y}}$, a prediction link yields a discrete prediction in $\widehat{\mathcal{Y}}$ from a score $\bm{\theta}\in\mathbb{R}^{\varrho}$ obtained through minimizing the surrogate loss. Thus a loss and link are the two sides of the same coin. Unlike the standard argmax link in multiclass classification [65, 68], we create an alternative argmax-like prediction link based on the minimizer of (10).
Definition 10 ($\bm{\pi}$-argmax link)
Let $\Pi:\mathbb{R}^{\varrho}\to 2^{\Delta^{N}}$ be a set-valued map defined as follows:
$\Pi(\bm{\theta})\coloneqq\mathop{\mathrm{missing}}{argmin}_{\bm{\pi}\in\Delta^{N}}\,\Omega^{*}(\bm{\theta}+\mathcal{L}^{\bm{\rho}}\bm{\pi})\quad\text{for any $\bm{\theta}\in\mathbb{R}^{\varrho}$}.$ (12)
Let $\bm{\pi}:\mathbb{R}^{\varrho}\to\Delta^{N}$ be a selector of $\Pi$ such that $\bm{\pi}(\bm{\theta})\in\Pi(\bm{\theta})$ for all $\bm{\theta}\in\mathbb{R}^{\varrho}$. Then the $\bm{\pi}$-argmax link $\varphi:\mathbb{R}^{\varrho}\to\widehat{\mathcal{Y}}$ is defined as
$\varphi(\bm{\theta})\in\mathop{\mathrm{missing}}{argmax}_{t\in\widehat{\mathcal{Y}}}\,\pi_{t}(\bm{\theta})\quad\text{for any $\bm{\theta}\in\mathbb{R}^{\varrho}$},$
where the tie can be broken arbitrarily.
The existence of such links is guaranteed as long as $\Pi(\bm{\theta})$ is non-empty for every $\bm{\theta}\in\mathbb{R}^{\varrho}$, which is guaranteed by Lemma 8. While the standard argmax link returns a prediction $t\in\widehat{\mathcal{Y}}$ with the maximum score $\theta_{t}$, the ${\bm{\pi}}$-argmax link returns $t$ with the maximum $\pi_{t}$. This probabilistic quantity $\bm{\pi}\in\Pi(\bm{\theta})$ defined via (10) distorts the original score $\bm{\theta}$ by leveraging the target loss structure $\mathcal{L}^{\bm{\rho}}$.
3.2 Convexity and Smoothness
Before discussing surrogate regret bounds, we verify that convolutional Fenchel–Young losses are indeed convex and smooth, implied by the conjugacy between smoothness and strong convexity [28].
Corollary 11 (Convexity and smoothness of $L_{\Omega_{T}}$)
Suppose that a discrete target loss $\ell$ enjoys $(\bm{\rho},\bm{\ell}^{\bm{\rho}})$-decomposition. For a base negentropy $\Omega$, we additionally suppose that Condition 1 is satisfied. If $\Omega$ is strictly convex on $\mathrm{dom}(\Omega)$, $L_{\Omega_{T}}(\cdot,y)$ is convex and differentiable over $\mathbb{R}^{\varrho}$ for any $y\in\mathcal{Y}$. If $\Omega$ is additionally strongly convex on $\mathrm{dom}(\Omega)$, $L_{\Omega_{T}}(\cdot,y)$ is smooth over $\mathbb{R}^{\varrho}$ for any $y\in\mathcal{Y}$.
There are several exemplar base entropies fulfilling the conditions of Theorem 11. For example, the squared norm $\Omega(\bm{p})=\frac{1}{2}\|\bm{p}\|^{2}$ is strongly convex with the self-conjugate $\Omega^{*}(\bm{\theta})=\frac{1}{2}\|\bm{\theta}\|^{2}$. In this case, $\Omega$ is effective over the entire $\mathbb{R}^{\varrho}$ and hence includes $\mathrm{conv}(\bm{\rho}(\mathcal{Y}))$, satisfying Condition 1. In light of the target-loss decomposition (3), we have $\mathrm{conv}(\bm{\rho}(\mathcal{Y}))=\Delta^{K}$. In this case, the Shannon negentropy $\Omega(\bm{p})=\langle\bm{p},\ln\bm{p}\rangle+\mathbb{I}_{\Delta^{K}}(\bm{p})$ is strongly convex on $\Delta^{K}$ with its conjugate $\Omega^{*}(\bm{\theta})=\ln\langle\exp(\bm{\theta}),\bm{1}\rangle$ satisfying $\mathrm{dom}(\Omega^{*})=\mathbb{R}^{K}$, where $\ln$ and $\exp$ are element-wise. Hence, both base negentropy functions yield convex and smooth $L_{\Omega_{T}}$.
While the surrogate loss $L_{\Omega_{T}}$ is now ensured to be convex and smooth, its gradient calculation, which is the basis for gradient-based optimization [11, 9], remains non-trivial. This is due to the gradient of the term $\min_{\bm{\pi}\in\Delta^{N}}\Omega^{*}(\bm{\theta}+\mathcal{L}^{\bm{\rho}}\bm{\pi})$, which cannot be written analytically in general. We show that its gradient calculation can be reduced to the calculation of $\Pi(\bm{\theta})$ through the following variant of envelope theorems.
Lemma 12 (Envelope theorem)
Suppose Condition 1 holds and $\Omega$ is strictly convex on $\mathrm{dom}(\Omega)$. For any $\bm{\theta}\in\mathbb{R}^{\varrho}$ and $\bm{\pi}\in\Pi(\bm{\theta})$, we have
$\nabla_{\bm{\theta}}\Bigl{[}\min_{\bm{\pi}\in\Delta^{N}}\Omega^{*}(\bm{\theta}+\mathcal{L}^{\bm{\rho}}\bm{\pi})\Bigr{]}=\nabla\Omega^{*}(\bm{\theta}+\mathcal{L}^{\bm{\rho}}\bm{\pi}).$ (13)
Deviating from the standard envelope theorems [9, 7], we carefully addresses the non-uniqueness of the minimizers. Given this, the gradient of the convolutional Fenchel–Young loss (11) is accessed via $\nabla_{\bm{\theta}}L_{\Omega_{T}}(\bm{\theta},y)=\nabla\Omega^{*}(\bm{\theta}+\mathcal{L}^{\bm{\rho}}\bm{\pi})-\bm{\rho}(y)$. At the core of its proof, the differentiability of $\Omega_{T}^{*}$ (indirectly obtained via Theorem 11) is vital.
3.3 Linear Surrogate Regret Bounds
Now we exhibit linear surrogate regret bounds with the convolutional Fenchel–Young loss.
Theorem 13 (Linear surrogate regret bound)
Consider a target loss with $(\bm{\rho},\bm{\ell}^{\bm{\rho}})$-decomposition. For an negentropy function $\Omega:\mathbb{R}^{\varrho}\to\overline{\mathbb{R}}$, suppose Condition 1 holds. For any $\bm{\pi}$-argmax link $\varphi$, $(L_{\Omega_{T}},\varphi)$ admits the following surrogate regret bound:
$\mathtt{Regret}_{\ell}(\varphi(\bm{\theta}),\bm{\eta})\leq N\mathtt{Regret}_{L_{\Omega_{T}}}\!(\bm{\theta},\bm{\eta}),~{}~{}~{}\text{for any~{}}(\bm{\theta},\bm{\eta})\in\mathbb{R}^{\varrho}\times\Delta^{K}.$ (14)
The regret bound (13) indeed has the linear rate $\psi(r)=Nr$ in the form (1), while $L_{\Omega_{T}}$ can be naturally made convex and smooth as discussed in Section 3.2. This is a remarkable consequence because many authors have previously implied the impossibility to overcome the barrier of the square-root regret rate with convex smooth surrogate losses [33, 53, 25, 4]. The crux of this success lies in the infimal convolution structure in (10), which enables us to additively decompose the conjugate $\Omega_{T}^{*}$, exploited below. The constant $N$ will be improved in Section 3.4.
From now on we sketch the regret bound proof. The following lemma is a cornerstone herein.
Lemma 14 (Lower bound of surrogate regret)
Assume the same set of conditions as in Theorem 13. For any $\bm{\theta}\in\mathbb{R}^{\varrho}$ and $\bm{\pi}\in\Pi(\bm{\theta})$, we have the following inequality:
$\sum_{t=1}^{N}\pi_{t}\mathtt{Regret}_{\ell}(t,\bm{\eta})\leq\mathtt{Regret}_{L_{\Omega_{T}}}\!(\bm{\theta},\bm{\eta}),~{}~{}~{}\text{for any~{}}(\bm{\theta},\bm{\eta})\in\mathbb{R}^{\varrho}\times\Delta^{K}.$ (15)
Its proof, formally given in Appendix A, hinges on the following additive regret decomposition:
$\mathtt{Regret}_{L_{\Omega_{T}}}\!(\bm{\theta},\bm{\eta})=\underbrace{R_{L_{\Omega}}(\bm{\theta}+\mathcal{L}^{\bm{\rho}}\bm{\pi},\bm{\eta})}_{\text{Risk of Fenchel--Young loss $L_{\Omega}$ $\geq 0$}}+\underbrace{\sum\nolimits_{t\in\widehat{\mathcal{Y}}}\pi_{t}\mathtt{Regret}_{\ell}(t,\bm{\eta})}_{\text{Convex combination of the target regret}}\!\forall(\bm{\theta},\bm{\pi})\in\mathbb{R}^{\varrho}\times\Pi(\bm{\theta}),$
which directly implies the lower bound (15) because the Fenchel–Young loss $L_{\Omega}$ is non-negative. Note that this additive decomposition is indispensable for the linear lower bound (15). This becomes possible just because we create the convolutional negentropy $\Omega_{T}$ based on the additive form in (8), which yields the additive form $\bm{\theta}+\mathcal{L}^{\bm{\rho}}\bm{\pi}$ in the conjugate expression (10). All of these are thanks to the structure of the infimal convolution. With this lemma, Theorem 13 can be proven immediately.
Proof of Theorem 13 Fix any $\bm{\theta}\in\mathbb{R}^{\varrho}$ and choose any $\bm{\pi}\in\Delta^{N}$ out of $\Pi(\bm{\theta})$. Then we have
$\forall\bm{\eta}\in\Delta^{K},\;\;\mathtt{Regret}_{L_{\Omega_{T}}}\!(\bm{\theta},\bm{\eta})$ $\geq\sum\nolimits_{t=1}^{N}\pi_{t}\mathtt{Regret}_{\ell}(t,\bm{\eta})$ (by (15))
$\geq\pi_{\varphi(\bm{\theta})}\mathtt{Regret}_{\ell}(\varphi(\bm{\theta}),\bm{\eta})$ (because $\pi_{t}\geq 0$ for any $t\in\widehat{\mathcal{Y}}$)
$\geq\mathtt{Regret}_{\ell}(\varphi(\bm{\theta}),\bm{\eta})/N.$ (by definition of $\varphi$ in Defintion 10)
Finally, we achieve the main result by combining Theorems 11 and 13. We constructively prove the existence of convex smooth linear-regret surrogate losses, by noting that both strongly convex $\Omega$ satisfying Condition 1 and $\bm{\pi}$-argmax link do exist.
Theorem 15 (Main result)
Consider a target loss $\ell$ with $(\bm{\rho},\bm{\ell}^{\bm{\rho}})$-decomposition. For an negentropy function $\Omega:\mathbb{R}^{\varrho}\to\overline{\mathbb{R}}$ satisfying Condition 1, suppose that $\Omega$ is strongly convex on $\mathrm{dom}(\Omega)$. Then the convolutional Fenchel–Young loss $L_{\Omega_{T}}(\cdot,y)$ is convex, differentiable, and smooth over $\mathbb{R}^{\varrho}$ for any $y\in\mathcal{Y}$. Moreover, with $\bm{\pi}$-argmax link $\varphi$, $(L_{\Omega_{T}},\varphi)$ enjoys the linear surrogate regret bound (14).
3.4 Improving Constant of Linear Surrogate Regret Bounds
The constant $N$ in the linear surrogate regret bound in (14) can be prohibitively large, leading to a vacuous bound. This issue is significant in structured prediction as studied in Osokin et al. [50]. For example, in multilabel classification with the Hamming loss, $N=2^{\varrho}$ is the number of all the potential binary label predictions that increases exponentially with $\varrho$, the number of binary labels. In top-$k$ classification, $N=\binom{K}{k}$ is the number of all possible size-$k$ subsets of the class space. Thankfully, the geometry property of the problem (10) provides a promising scheme for reducing the dependency on prediction number $N$, which induces the following improved link.
Corollary 16 (Improved surrogate regret bound)
Suppose Condition 1 holds. There exists $\tilde{\bm{\pi}}:\mathbb{R}^{\varrho}\to\Delta^{N}$ such that $\tilde{\bm{\pi}}(\bm{\theta})\in\Pi(\bm{\theta})$ and $\|\tilde{\bm{\pi}}(\bm{\theta})\|_{0}\leq\mathrm{affdim}(\mathcal{L}^{\bm{\rho}})+1$ for any $\bm{\theta}\in\mathbb{R}^{\varrho}$. Moreover, with the induced $\tilde{\bm{\pi}}$-argmax link $\varphi_{\tilde{\bm{\pi}}}:\mathbb{R}^{\varrho}\to\Delta^{N}$, $(L_{\Omega_{T}},\varphi_{\tilde{\bm{\pi}}})$ admits the following surrogate regret bound:
$\mathtt{Regret}_{\ell}(\varphi_{\tilde{\bm{\pi}}}(\bm{\theta}),\bm{\eta})\leq[\mathrm{affdim}(\mathcal{L}^{\bm{\rho}})+1]\,\mathtt{Regret}_{L_{\Omega_{T}}}\!(\bm{\theta},\bm{\eta}),\quad\text{for any~{}}(\bm{\theta},\bm{\eta})\in\mathbb{R}^{\varrho}\times\Delta^{K},$ (16)
where $\mathrm{affdim}(\mathcal{L}^{\bm{\rho}})$ is the dimension of the affine hull of the column vectors of $\mathcal{L}^{\bm{\rho}}$.
Proof.  Fix an arbitrary $\bm{\theta}\in\mathbb{R}^{\varrho}$. Since $\Pi(\bm{\theta})$ is non-empty, we can select $\bm{\pi}$ from $\Pi(\bm{\theta})$. Note that $\mathcal{L}^{\bm{\rho}}\bm{\pi}\in\mathrm{conv}(\{\bm{\ell}^{\bm{\rho}}(t)\}_{t\in\widehat{\mathcal{Y}}})$. According to Carathéodory’s theorem, there exists $\tilde{\bm{\pi}}$ such that $\|\tilde{\bm{\pi}}\|_{0}\leq\mathrm{affdim}(\mathcal{L}^{\bm{\rho}})+1$ and $\mathcal{L}^{\bm{\rho}}\tilde{\bm{\pi}}=\mathcal{L}^{\bm{\rho}}\bm{\pi}$ is also in $\Pi(\bm{\theta})$. Since $\bm{\theta}$ is chosen arbitrarily, we can then construct a single-valued function $\tilde{\bm{\pi}}:\mathbb{R}^{\varrho}\to\Delta^{N}$ such that $\tilde{\bm{\pi}}(\bm{\theta})\in\Pi(\bm{\theta})$ and $\|\tilde{\bm{\pi}}(\bm{\theta})\|_{0}\leq\mathrm{affdim}(\mathcal{L}^{\bm{\rho}})+1$. Noting that $\max_{t\in\widehat{\mathcal{Y}}}\tilde{\pi}_{t}(\bm{\theta})>{1}/[\mathrm{affdim}(\mathcal{L}^{\bm{\rho}})+1]$, we can complete the rest of the proof similarly to Theorem 13.
Note that we have $\mathrm{affdim}(\mathcal{L}^{\bm{\rho}})\leq\min\{N,\varrho\}$, which indicates that the dependency on $N$ can be largely reduced when $\varrho\ll N$. For example, in top-$k$ classification, $\bm{\rho}(y)=\bm{e}_{y}$ is used with $\varrho=K$, which is much smaller than $N=\binom{K}{k}$. In multilabel classification with the Hamming loss, $\varrho=\ln_{2}N$ is logarithmically smaller than $N$ when the binary encoding $\bm{\rho}(y)=\bm{\nu}(y)$ is used (see Section 2.1).
3.5 Bonus: Fisher-consistent Probability Estimator
In this subsection, we discuss a benefit of convex smooth surrogate losses beyond discrete prediction problems. Oftentimes people are interested in a probability estimator over possible prediction outcomes, recovering from surrogate risk minimizers, as in selective classification [16]. Herein non-smooth surrogate losses have been unfavored because of lacking reasonable probability estimators [42, 15]. Ramaswamy et al. [53] conjectures the impossibility to establish a probability estimator and linear regret bound simultaneously. By contrast, we give a Fisher-consistent probability estimator of the linear property $\mathbb{E}_{y\sim\bm{\eta}}[\bm{\rho}(y)]$ under the convolutional Fenchel–Young loss, without sacrificing the fast regret bound.
Theorem 17 (Fisher-consistent probability estimator)
Suppose Condition 1 holds and $\Omega$ is strictly convex on $\mathrm{dom}(\Omega)$. For any $\bm{\eta}\in\mathrm{relint}(\Delta^{K})$, the surrogate risk $R_{L_{\Omega_{T}}}(\cdot,\bm{\eta})$ is minimized at $\bm{\theta}^{*}\in\mathbb{R}^{\varrho}$ such that $\nabla\Omega^{*}(\bm{\theta}^{*}+\mathcal{L}^{\bm{\rho}}\bm{\pi})=\mathbb{E}_{y\sim\bm{\eta}}[\bm{\rho}(y)]$ for any $\bm{\pi}\in\Pi(\bm{\theta}^{*})$.
According to the result above, we can first solve $\Pi(\bm{\theta})$ in (12) and select $\bm{\pi}\in\Pi(\bm{\theta})$ to use $\nabla\Omega^{*}(\bm{\theta}+\mathcal{L}^{\bm{\rho}}\bm{\pi})$ as an estimate of $\mathbb{E}_{y\sim\bm{\eta}}[\bm{\rho}(y)]$ with consistency guarantee. When we follow the trivial loss decomposition (3) with $\bm{\rho}(y)=\bm{e}_{y}$, the estimand is $\mathbb{E}_{y\sim\bm{\eta}}[\bm{\rho}(y)]=\bm{\eta}\in\Delta^{K}$. For this reason, we say $\nabla\Omega^{*}(\bm{\theta}^{*}+\mathcal{L}^{\bm{\rho}}\bm{\pi})$ is a “probability” estimator. For example, let us recap the encoding of multilabel classification in Section 2.1, where $y\in\mathcal{Y}$ is a multilabel among all $2^{\varrho}$ possible combinations of $\varrho$ binary labels. A multilabel $y$ is encoded by $\bm{\nu}(y)\in\{0,1\}^{\varrho}$, and the Hamming loss is decomposed with $\bm{\rho}(y)=1-2\bm{\nu}(y)$. Here $\mathbb{E}_{y\sim\bm{\eta}}[\bm{\nu}(y)]$ reads $\bar{\bm{\nu}}=[\mathrm{Prob}(\bm{\nu}_{i}(y)=1)]_{i=1}^{\varrho}$, whose $i$-th element indicates the likelihood of the binary class $i$ being positive. Through the relation $\bm{\rho}=1-2\bm{\nu}$, the probability estimator $\nabla\Omega^{*}(\bm{\theta}+\mathcal{L}^{\bm{\rho}}\bm{\pi})$ is capable of recovering $\bar{\bm{\nu}}$ by $[1-\nabla\Omega^{*}(\bm{\theta}+\mathcal{L}^{\bm{\rho}}\bm{\pi})]/2$.
4 Example: Multiclass Classification
In multiclass classification, the class space and the prediction space are the same, that is, $\mathcal{Y}=\widehat{\mathcal{Y}}=[K]$. For an input instance with multiclass probability $\bm{\eta}\in\Delta^{K}$, the goal is to predict the most likely class label $t\in\mathrm{argmax}_{t\in\mathcal{Y}}\,\eta_{t}$. The standard target loss is the 0-1 loss $\ell_{01}(t,y)=[\![t\not=y]\!]$.
Firstly, we adopt the decomposition (3) for this task, that is, $\bm{\ell}^{\bm{\rho}}(t)=[\ell_{01}(t,1),\cdots,\ell_{01}(t,K)]=\bm{1}-\bm{e}_{t}$ and $\bm{\rho}(y)=\bm{e}_{y}$, which corresponds to the one-hot encoding commonly used in this task. In this case, $N=\varrho=K$, and $\mathcal{L}^{\bm{\rho}}=\bm{1}\bm{1}^{\top}-I$, where $I$ is the identity matrix.
Next, we move on to the convolutional negentropy $\Omega_{T}$. For the choice of $\Omega$, we use the Shannon negentropy, which induces the celebrated cross-entropy loss in the original Fenchel–Young loss framework [10, Table 1]. Its conjugate is the log-sum-exp function $\Omega^{*}(\bm{\theta})=\ln\langle\exp(\bm{\theta}),\bm{1}\rangle$. Then we can calculate the conjugate of $\Omega_{T}$ based on Lemma 8:
$\Omega_{T}^{*}(\bm{\theta})=\min_{\bm{\pi}\in\Delta^{K}}\Omega^{*}(\bm{\theta}+\mathcal{L}^{\bm{\rho}}\bm{\pi})=\min_{\bm{\pi}\in\Delta^{K}}\ln\left\langle\exp(\bm{\theta}+\bm{1}-\bm{\pi}),\bm{1}\right\rangle.$ (17)
Since $\Omega$ is proper convex and l.s.c. with $\mathrm{conv}(\bm{\rho}(\mathcal{Y}))=\Delta^{K}=\mathrm{dom}(\Omega)$ and $\mathrm{dom}(\Omega^{*})=\mathbb{R}^{K}$, $\Omega$ satisfies Condition 1. Furthermore, since $\Omega$ is strongly convex on $\Delta^{K}$, we can finally derive the convolutional Fenchel–Young loss (11) by nothing $\Omega_{T}(\bm{\rho}(y))=0$ for all $y\in\mathcal{Y}$, as follows:
$L_{\Omega_{T}}(\bm{\theta},y)=\Omega_{T}^{*}(\bm{\theta})+\Omega_{T}(\bm{\rho}(y))-\langle\bm{\theta},\bm{\rho}(y)\rangle=\min_{\bm{\pi}\in\Delta^{K}}\ln\left\langle\exp(\bm{\theta}+\bm{1}-\bm{\pi}),\bm{1}\right\rangle-\theta_{i},$ (18)
which is ensured to be convex and smooth in $\bm{\theta}$ thanks to Theorem 11. While sharing the form of cross-entropy loss $L_{\Omega}(\bm{\theta},y)=\ln\langle\exp(\bm{\theta}),\bm{1}\rangle-\theta_{y}$, it further incorporates an additional bounded perturbation.
Algorithm 1 Exact Solution of (17) in $\mathcal{O}(K\ln K)$
1:  Sort $\bm{\theta}\in\mathbb{R}^{K}$ such that $\theta_{[1]}\geq\cdots\geq\theta_{[K]}$.
2:  $n\leftarrow\max\{k\in[K]:~{}1+k\theta_{[k]}>\sum_{i=1}^{k}\theta_{[i]}\}$.
3:  $\tau(\bm{\theta})\leftarrow\frac{\sum_{i=1}^{n}\theta_{[i]}-1}{n}$.
4:  $\pi_{\log}(\theta)_{i}\leftarrow\max\{\theta_{i}-\tau(\bm{\theta}),0\}$.
Solving the minimization problem (17) is important from the computational aspects, including the gradient calculation of $L_{\Omega_{T}}$ and accessing to the probability estimator given by Theorem 17. We provide Algorithm 1 to solve (17) with $\mathcal{O}(K\ln K)$ time.
Corollary 18
For any $\bm{\theta}\in\mathbb{R}^{K}$, the problem (17) has a unique minimizer $\bm{\pi}_{\log}(\bm{\theta})\in\Pi(\bm{\theta})$, which can be obtained in $\mathcal{O}(K\ln K)$ time by Algorithm 1.
The proof is deferred to Appendix B.1, following from the KKT conditions. Eventually we have
$\nabla\Omega_{T}^{*}(\bm{\theta})=\mathrm{softmax}\left(\bm{\theta}+\bm{1}-\bm{\pi}_{\log}(\bm{\theta})\right),\;\;\text{where~{}~{}}\mathrm{softmax}(\bm{\theta})_{y}=\frac{\exp(\theta_{y})}{\sum_{i=1}^{K}\exp(\theta_{i})}\;\;\text{for $y\in[K]$}.$
Algorithm 1 comes with a significant resemblance with the sparsemax computation [41, Algorithm 1], which is determined by the similar structure shared by (17) and Euclidean projection problem [41, (2)]. Since $\Pi(\bm{\theta})$ is a singleton, we have the unique $\bm{\pi}_{\log}$-argmax link $\varphi_{\bm{\pi}_{\log}}$ (Definition 10). While we need to solve the problem (17) for every $\bm{\theta}$ to have access to $\nabla_{\bm{\theta}}L_{\Omega_{T}}(\bm{\theta},y)$, the test-time prediction by the $\bm{\pi}_{\log}$-argmax link is much cheaper, without requiring Algorithm 1. The proof can be found in Appendix B.2.
Proposition 19
A prediction link $\varphi$ is the $\bm{\pi}_{\log}$-argmax link if and only if $\varphi(\bm{\theta})\in\mathrm{argmax}_{t\in\mathcal{Y}}\theta_{t}$.
According to this result, we can simply use the class label with the largest score in test time.
Remark 20
Under classification, a surrogate loss with a regret bound is classification-calibrated, which indicates that the surrogate risk minimization eventually leads to the Bayes-optimal classifier. In literature, Blondel [8] and Wang and Scott [68] have investigated sufficient conditions for Fenchel–Young losses to be classification-calibrated. Therein the base negentropy is assumed to be of Legendre-type or twice differentiable. Our convolutional Fenchel–Young losses are interesting because we do not require these conditions to yield both the smoothness and a regret bound. Thus it remains open to relax these existing sufficient conditions for Fenchel–Young losses further.
5 Discussions
Efficient gradient calculation.
In general discrete prediction problems, we need to solve the minimization (10) to take a gradient of $L_{\Omega_{T}}$, which is potentially demanding over a high-dimensional domain $\Delta^{N}$. To have access to the gradient, we can alternatively solve
$\min_{\bm{\nu}\in V}\Omega^{*}(\bm{\theta}+\bm{\nu}),\quad\text{where~{}~{}}V\coloneqq\mathrm{conv}\Bigl{(}\{\bm{\ell}^{\bm{\rho}}(t)\}_{t\in[N]}\Bigr{)}.$ (19)
To see this, let us denote the minimizer of (19) by $\bm{\nu}^{*}$. Then (19) is an equivalent optimization problem to (10) because there exists $\bm{\pi}\in\Pi(\bm{\theta})$ such that $\bm{\nu}^{*}=\mathcal{L}^{\bm{\rho}}\bm{\pi}$. Eventually $\nabla\Omega^{*}(\bm{\theta}+\bm{\nu}^{*})$ serves as an alternative to the gradient formula (13). Thus we can reduce the dimensionality of the optimization problem. For example, in multilabel classification (Section 2.1), $V=[0,1]^{\varrho}$ has a logarithmically smaller optimization dimensionality than $N=2^{\varrho}$. We can solve (19) with this box constraint efficiently by using the standard L-BFGS solver [30].
Randomized prediction link.
Recall that the $\bm{\pi}$-argmax link $\varphi$ (Definition 10) deterministically outputs in the probability simplex $\Delta^{N}$. Instead we can define a randomized link $\tilde{\varphi}$ such that $\Pr[\tilde{\varphi}(\bm{\theta})=t]=\pi_{t}(\bm{\theta})$. This yields a better regret bound by Lemma 14, as follows:
$\mathbb{E}_{\tilde{\varphi}(\bm{\theta})}[\mathtt{Regret}_{\ell}(\tilde{\varphi}(\bm{\theta}),\bm{\eta})]=\sum_{t=1}^{N}\pi_{t}(\bm{\theta})\,\mathtt{Regret}_{\ell}(t,\bm{\eta})\leq\mathtt{Regret}_{L_{\Omega_{T}}}\!(\bm{\theta},\bm{\eta})\quad\forall(\bm{\theta},\bm{\eta})\in\mathbb{R}^{\varrho}\times\Delta^{K},$
where the expectation is taken over the randomness of $\tilde{\varphi}$. Although we adopt the expected target regret differently from Theorem 13, the constant of regret bounds is strikingly improved from $N$ to $1$, which is dimension-free. Sakaue et al. [59] observes a similar regret improvement by a randomized link for online structured prediction with Fenchel–Young losses.
Strong convexity.
Since the convolutional Fenchel–Young loss possesses both smoothness and a link with a linear surrogate regret bound, it cannot be strongly convex. This can be obtained as an immediate corollary of Frongillo and Waggoner [25, Theorem 4], which proves the impossibility of the combination of smoothness, strong convexity, and a linear surrogate regret bound.333 Strictly speaking, Frongillo and Waggoner [25] assumes a target loss is non-redundant and a surrogate loss is smooth and locally strongly convex at a decision boundary.
This incompatibility can also be observed from the construction of the convolutional Fenchel–Young losses: $\Omega_{T}$ is non-smooth under non-redundant $\ell$ due to the piecewise linear $\min_{t}\left\langle\bm{p},\bm{\ell}^{\bm{\rho}}(t)\right\rangle$ and thus its conjugate is not strongly convex due to the duality between smoothness and strong convexity [28]. To put it differently, $L_{\Omega_{T}}$ cannot be strongly convex.
Acknoledgement
YC is supported by Google PhD Fellowship program. HB is supported by JST PRESTO (Grant No. JPMJPR24K6), Japan.
Appendix A Deferred Proofs in Section 3
For a clear representation in this section, we denote by $T(\bm{p})$ the polyhedral convex function $-\min_{t\in\widehat{\mathcal{Y}}}\left\langle\bm{p},\bm{\ell}^{\bm{\rho}}(t)\right\rangle$, which is the negative of the affinely transformed target Bayes risk in (9). Here, $T$ is proper convex and l.s.c., and we have $\Omega_{T}=\Omega+T$.
A.1 Proof of Lemma 8
Proof.  Since $\Omega$ and $T$ are proper convex and l.s.c., so are $\Omega_{T}$ and thus $\Omega_{T}^{*}$. Then we prove (10) using the infimal convolution. First, by noting that $T$ is nothing else but the support function of the closed convex set $\mathrm{conv}(\{-\bm{\ell}^{\bm{\rho}}(t)\}_{t\in\widehat{\mathcal{Y}}})$, we can express $T^{*}$ as follows:
$T^{*}(\bm{\theta})=\sup_{\bm{p}\in\mathbb{R}^{\varrho}}\left[\langle\bm{\theta},\bm{p}\rangle-\max_{t\in\widehat{\mathcal{Y}}}\left\langle\bm{p},-\bm{\ell}^{\bm{\rho}}(t)\right\rangle\right]=\mathbb{I}_{\mathrm{conv}\left(\{-\bm{\ell}^{\bm{\rho}}(t)\}_{t\in\widehat{\mathcal{Y}}}\right)}(\bm{\theta}),$
where we use the conjugacy relationship between a support function and indicator function of a closed convex set $\mathrm{conv}(\{-\bm{\ell}^{\bm{\rho}}(t)\}_{t\in\widehat{\mathcal{Y}}})$ [57, Section 13]. According to Condition 1 and the definition of $T$, we see that both $\Omega$ and $T$ are proper convex and $T$ is continuous on $\mathrm{dom}(\Omega)$. Then we use the infimal convolution [13] to derive the conjugate of $\Omega_{T}$:
$\Omega_{T}^{*}(\bm{\theta})$ $=\inf_{\bm{\theta}^{\prime}\in\mathbb{R}^{d}}\bigg{[}\,\Omega^{*}(\bm{\theta}-\bm{\theta}^{\prime})+\mathbb{I}_{\mathrm{conv}\left(\{-\bm{\ell}^{\bm{\rho}}(t)\}_{t\in\widehat{\mathcal{Y}}}\right)}(\bm{\theta}^{\prime})\,\bigg{]}$
$=\inf_{\bm{\theta}^{\prime}\in\mathrm{conv}\left(\{-\bm{\ell}^{\bm{\rho}}(t)\}_{t\in\widehat{\mathcal{Y}}}\right)}\Omega^{*}(\bm{\theta}-\bm{\theta}^{\prime})$
$=\inf_{\bm{\pi}\in\Delta^{N}}\Omega^{*}(\bm{\theta}+\mathcal{L}^{\bm{\rho}}\bm{\pi}).$
Next we show that the infimum is indeed achieved by some $\bm{\pi}\in\Delta^{N}$. Note that for any $\bm{\theta}\in\mathbb{R}^{\varrho}$, the set $\Theta\coloneqq\{\bm{\theta}+\mathcal{L}^{\bm{\rho}}\bm{\pi}:\bm{\pi}\in\Delta^{N}\}$ is compact and non-empty, and that $\Omega^{*}$ is l.s.c. on $\Theta$ since $\mathrm{dom}({\Omega^{*}})=\mathbb{R}^{\varrho}$. Then the infimum of $\Omega^{*}(\bm{\theta})$ is achieved by some $\bm{\theta}^{\prime\prime}\in\Theta$, and there must exist $\bm{\pi}\in\Delta^{N}$ such that $\bm{\theta}+\mathcal{L}^{\bm{\rho}}\bm{\pi}=\bm{\theta}^{\prime\prime}$ by the definition of $\Theta$, which complete the proof of the existence of the minimizer.
Finally we see that for any $\bm{\theta}\in\mathbb{R}^{\varrho}$, there exists $\bm{\pi}\in\Delta^{N}$ such that $\Omega_{T}^{*}(\bm{\theta})=\Omega^{*}(\bm{\theta}+\mathcal{L}^{\bm{\rho}}\bm{\pi})$, which is smaller than +$\infty$+ ∞ since $\mathrm{dom}(\Omega^{*})=\mathbb{R}^{\varrho}$ and $\bm{\theta}+\mathcal{L}^{\bm{\rho}}\bm{\pi}\in\mathbb{R}^{\varrho}$. This indicates that $\mathrm{dom}(\Omega_{T}^{*})=\mathbb{R}^{\varrho}$ because $\bm{\theta}$ is chosen arbitrarily.
A.2 Proof of Theorem 11
Proof.  In the convolutional Fenchel–Young loss (11), the term $\Omega_{T}(\bm{\rho}(y))-\langle\bm{\theta},\bm{\rho}(y)\rangle$ is linear. Hence it suffices to prove the convexity and smoothness of the conjugate $\min_{\bm{\pi}\in\Delta^{N}}\Omega^{*}(\bm{\theta}+\mathcal{L}^{\bm{\rho}}\bm{\pi})$.
When $\Omega$ is strictly convex, $\Omega_{T}$ is also strictly convex since $T$ is convex. Since the strict convexity holds on $\mathrm{dom}(\Omega_{T})$, $\Omega_{T}$ is further essentially strictly convex, that is, strictly convex on every convex subset of $\mathrm{dom}(\partial\Omega_{T})$ [57, p253]. According to Rockafellar [57, Theorem 26.3], $\Omega_{T}^{*}$ is essentially smooth, that is, differentiable throughout non-empty $\mathrm{int}(\mathrm{dom}(\Omega_{T}^{*}))$ with $\|\nabla\Omega_{T}^{*}(\bm{\theta})\|$ diverging to +$\infty$+ ∞ when $\bm{\theta}$ approaches a boundary point of $\mathrm{int}(\mathrm{dom}(\Omega_{T}^{*}))$, which immediately indicates the differentiability on $\mathbb{R}^{\varrho}$.
[…]
A.3 Proof of Lemma 12
Proof.  According to Condition 1, the strict convexity of $\Omega$, and the proof of Theorem 11, both $\Omega_{T}^{*}$ and $\Omega^{*}$ are differentiable on $\mathbb{R}^{\varrho}$. In addition, we have $\mathrm{dom}(\Omega)=\mathrm{dom}(\Omega_{T})$ according to (8).
Note that $\Omega^{*}(\bm{\theta}+\mathcal{L}^{\bm{\rho}}\bm{\pi})$ is convex in $\bm{\pi}$. Then its first-order optimal condition for any $\bm{\pi}\in\Pi(\bm{\theta})$ reads
$\bigl{\langle}\nabla\Omega^{*}(\bm{\theta}+\mathcal{L}^{\bm{\rho}}\bm{\pi}),\mathcal{L}^{\bm{\rho}}(\bm{\pi}^{\prime}-\bm{\pi})\bigr{\rangle}\geq 0\qquad\text{any $\bm{\pi}^{\prime}\in\Delta^{N}$.}$
First, for any $\bm{\pi}\in\Pi(\bm{\theta})$, we choose
$t\in\mathrm{argmin}_{t^{\prime}\in\widehat{\mathcal{Y}}}\langle\nabla\Omega^{*}(\bm{\theta}+\mathcal{L}^{\bm{\rho}}\bm{\pi}),\bm{\ell}^{\bm{\rho}}(t^{\prime})\rangle,$
then we have
$\Omega^{*}(\bm{\theta}+\mathcal{L}^{\bm{\rho}}\bm{\pi})$ =(A)${=}(\bm{\theta}+\mathcal{L}^{\bm{\rho}}\bm{\pi})^{\top}\nabla\Omega^{*}(\bm{\theta}+\mathcal{L}^{\bm{\rho}}\bm{\pi})-\Omega(\nabla\Omega^{*}(\bm{\theta}+\mathcal{L}^{\bm{\rho}}\bm{\pi}))$
$=\bm{\theta}^{\top}\nabla\Omega^{*}(\bm{\theta}+\mathcal{L}^{\bm{\rho}}\bm{\pi})+\left\langle\nabla\Omega^{*}(\bm{\theta}+\mathcal{L}^{\bm{\rho}}\bm{\pi}),\mathcal{L}^{\bm{\rho}}\bm{\pi}\right\rangle-\Omega(\nabla\Omega^{*}(\bm{\theta}+\mathcal{L}^{\bm{\rho}}\bm{\pi}))$
(A)${\leq}\bm{\theta}^{\top}\nabla\Omega^{*}(\bm{\theta}+\mathcal{L}^{\bm{\rho}}\bm{\pi})+\left\langle\nabla\Omega^{*}(\bm{\theta}+\mathcal{L}^{\bm{\rho}}\bm{\pi}),\mathcal{L}^{\bm{\rho}}\bm{e}_{t}\right\rangle-\Omega(\nabla\Omega^{*}(\bm{\theta}+\mathcal{L}^{\bm{\rho}}\bm{\pi}))$
$=\bm{\theta}^{\top}\nabla\Omega^{*}(\bm{\theta}+\mathcal{L}^{\bm{\rho}}\bm{\pi})+\langle\nabla\Omega^{*}(\bm{\theta}+\mathcal{L}^{\bm{\rho}}\bm{\pi}),\bm{\ell}^{\bm{\rho}}(t)\rangle-\Omega(\nabla\Omega^{*}(\bm{\theta}+\mathcal{L}^{\bm{\rho}}\bm{\pi}))$
=(B)${=}\bm{\theta}^{\top}\nabla\Omega^{*}(\bm{\theta}+\mathcal{L}^{\bm{\rho}}\bm{\pi})+\underbrace{\min_{t^{\prime}\in\widehat{\mathcal{Y}}}\langle\nabla\Omega^{*}(\bm{\theta}+\mathcal{L}^{\bm{\rho}}\bm{\pi}),\bm{\ell}^{\bm{\rho}}(t^{\prime})\rangle}_{=-T\left(\nabla\Omega^{*}(\bm{\theta}+\mathcal{L}^{\bm{\rho}}\bm{\pi})\right)}-\Omega(\nabla\Omega^{*}(\bm{\theta}+\mathcal{L}^{\bm{\rho}}\bm{\pi}))$
=(C)${=}\bm{\theta}^{\top}\nabla\Omega^{*}(\bm{\theta}+\mathcal{L}^{\bm{\rho}}\bm{\pi})-\Omega_{T}(\nabla\Omega^{*}(\bm{\theta}+\mathcal{L}^{\bm{\rho}}\bm{\pi}))$
(D)${\leq}\sup_{\bm{p}\in\mathrm{dom}(\Omega_{T})}\bm{\theta}^{\top}\bm{p}-\Omega_{T}(\bm{p})$
$=\Omega_{T}^{*}(\bm{\theta}),$
where (A) owes to the optimality of $\bm{\pi}\in\Pi(\bm{\theta})$, (B) holds by the definition of $t$, and (C) holds by the definition of the convolutional negentropy $\Omega_{T}$ (8). Since $\bm{\pi}\in\Pi(\bm{\theta})$, we have $\Omega^{*}(\bm{\theta}+\mathcal{L}^{\bm{\rho}}\bm{\pi})=\Omega_{T}^{*}(\bm{\theta})$ according to Lemma 8. Thus the above inequality is indeed an identity. In particular, (D) becomes an identity, which implies that the supremum of $\sup_{\bm{p}\in\mathrm{dom}(\Omega_{T})}\bm{\theta}^{\top}\bm{p}-\Omega_{T}(\bm{p})$ is achieved at $\nabla\Omega^{*}(\bm{\theta}+\mathcal{L}^{\bm{\rho}}\bm{\pi})$. Since the supremum is also achieved at $\nabla\Omega_{T}^{*}(\bm{\theta})$ and the maximizer is unique according to Rockafellar and Wets [58, Proposition 11.3] and the differentiability of $\Omega^{*}_{T}$, we have
$\nabla\Omega_{T}^{*}(\bm{\theta})=\nabla\Omega^{*}(\bm{\theta}+\mathcal{L}^{\bm{\rho}}\bm{\pi}).$
Since $\bm{\pi}\in\Pi(\bm{\theta})$ and $\bm{\theta}\in\mathbb{R}^{\varrho}$ are chosen arbitrarily, this concludes the proof.
A.4 Proof of Lemma 14
Proof.  First, we derive the Bayes risk of the convolutional Fenchel–Young loss $L_{\Omega_{T}}$ as follows:
$\underline{R}_{L_{\Omega_{T}}}(\bm{\eta})$ $=\inf_{\bm{\theta}\in\mathbb{R}^{\varrho}}R_{L_{\Omega_{T}}}(\bm{\theta},\bm{\eta})$
$=\inf_{\bm{\theta\in\mathbb{R}^{\varrho}}}\Bigl{[}\Omega_{T}^{*}(\bm{\theta})-\langle\bm{\theta},\mathbb{E}_{y\sim\bm{\eta}}[\bm{\rho}(y)]\rangle\Bigr{]}+\mathbb{E}_{y\sim\bm{\eta}}[\Omega_{T}(\bm{\rho}(y))]$
$=-\sup_{\bm{\theta\in\mathbb{R}^{\varrho}}}\Bigl{[}\langle\bm{\theta},\mathbb{E}_{y\sim\bm{\eta}}[\bm{\rho}(y)]\rangle-\Omega_{T}^{*}(\bm{\theta})\Bigr{]}+\mathbb{E}_{y\sim\bm{\eta}}[\Omega_{T}(\bm{\rho}(y))]$
$=-\Omega_{T}^{**}\Bigl{(}\mathbb{E}_{y\sim\bm{\eta}}[\bm{\rho}(y)]\Bigr{)}+\mathbb{E}_{y\sim\bm{\eta}}[\Omega_{T}(\bm{\rho}(y))]$
$=\mathbb{E}_{y\sim\bm{\eta}}[\Omega_{T}(\bm{\rho}(y))]-\Omega_{T}\Bigl{(}\mathbb{E}_{y\sim\bm{\eta}}[\bm{\rho}(y)]\Bigr{)}$
where the Fenchel–Moreau theorem is applied to proper convex and l.s.c. $\Omega_{T}^{*}$ (by Lemma 8) at the last identity. Then we can get the following regret lower bound for any $\bm{\pi}\in\Pi(\bm{\theta})$:
$\mathtt{Regret}_{L_{\Omega_{T}}}(\bm{\theta},\bm{\eta})$ $=R_{L_{\Omega_{T}}}(\bm{\theta},\bm{\eta})-\underline{R}_{L_{\Omega_{T}}}(\bm{\eta})$
$=\Omega_{T}^{*}(\bm{\theta})-\Bigl{\langle}\bm{\theta},\mathbb{E}_{y\sim\bm{\eta}}[\bm{\rho}(y)]\Bigr{\rangle}+\Omega_{T}\Bigl{(}\mathbb{E}_{y\sim\bm{\eta}}[\bm{\rho}(y)]\Bigr{)}$
$=\Omega^{*}(\bm{\theta}+\mathcal{L}^{\bm{\rho}}\bm{\pi})-\Bigl{\langle}\bm{\theta},\mathbb{E}_{y\sim\bm{\eta}}[\bm{\rho}(y)]\Bigr{\rangle}+\Omega\Bigl{(}\mathbb{E}_{y\sim\bm{\eta}}[\bm{\rho}(y)]\Bigr{)}+T\Bigl{(}\mathbb{E}_{y\sim\bm{\eta}}[\bm{\rho}(y)]\Bigr{)}$
=(A)${=}\underbrace{\Omega^{*}(\bm{\theta}+\mathcal{L}^{\bm{\rho}}\bm{\pi})-\Bigl{\langle}\bm{\theta}+\mathcal{L}^{\bm{\rho}}\bm{\pi},\mathbb{E}_{y\sim\bm{\eta}}[\bm{\rho}(y)]\Bigr{\rangle}+\Omega\Bigl{(}\mathbb{E}_{y\sim\bm{\eta}}[\bm{\rho}(y)]\Bigr{)}}_{\geq 0\quad\text{due to the Fenchel--Young inequality}}$
$~{}~{}~{}+\Bigl{\langle}\mathcal{L}^{\bm{\rho}}\bm{\pi},\mathbb{E}_{y\sim\bm{\eta}}[\bm{\rho}(y)]\Bigr{\rangle}+T\Bigl{(}\mathbb{E}_{y\sim\bm{\eta}}[\bm{\rho}(y)]\Bigr{)}$
(B)${\geq}\sum_{t=1}^{N}\pi_{t}\Bigl{\langle}\mathbb{E}_{y\sim\bm{\eta}}[\bm{\rho}(y)],\bm{\ell}^{\bm{\rho}}(t)\Bigr{\rangle}-\min_{t\in\widehat{\mathcal{Y}}}\Bigl{\langle}\mathbb{E}_{y\sim\bm{\eta}}[\bm{\rho}(y)],\bm{\ell}^{\bm{\rho}}(t)\Bigr{\rangle}$
=(C)${=}\sum_{t=1}^{N}\pi_{t}\mathtt{Regret}_{\ell}(t,\bm{\eta}),$
where (A) holds according to the explicit form of $\Omega_{T}^{*}$ in (10) and the definition of $\Pi(\bm{\theta})$, (B) owes to Fenchel–Young inequality, and (C) holds according to the definition of $(\bm{\rho},\bm{\ell}^{\bm{\rho}})$-decomposition in (4). This concludes the proof.
A.5 Proof of Theorem 17
Proof.  First of all, we prove that $\mathrm{relint}(\mathrm{conv}(\bm{\rho}(\mathcal{Y})))$ is in the image of $\nabla\Omega_{T}^{*}$ by contradiction. According to Rockafellar [57, Theorem 23.4], we have that $\partial\Omega_{T}(\bm{p})$ is non-empty for all $\bm{p}\in\mathrm{relint}(\mathrm{dom}(\Omega_{T}))$. Now suppose there exists $\bm{p}\in\mathrm{relint}(\mathrm{dom}(\Omega_{T}))$ such that $\bm{p}\not=\nabla\Omega_{T}^{*}(\bm{\theta})$ for any $\bm{\theta}\in\mathbb{R}^{\varrho}$. Since $\partial\Omega_{T}(\bm{p})$ is non-empty, there exists $\bm{\theta}^{\prime}\in\mathbb{R}^{\varrho}$ such that $\bm{\theta}^{\prime}=\nabla\Omega_{T}(\bm{p})$. By Rockafellar and Wets [58, Proposition 11.3] (applied on proper l.s.c. $\Omega_{T}$), we have $\bm{p}=\nabla\Omega_{T}^{*}(\bm{\theta^{\prime}})$, which contradicts the assumption $\bm{p}\not=\nabla\Omega_{T}^{*}(\bm{\theta}^{\prime})$. Thus we have verified that $\mathrm{relint}(\mathrm{dom}(\Omega_{T}))$ is in the image of $\nabla\Omega_{T}^{*}$. This additionally implies that $\mathrm{relint}(\mathrm{conv}(\bm{\rho}(\mathcal{Y})))\subseteq\mathrm{relint}(\mathrm{dom}(\Omega_{T}))$ is also in the image of $\nabla\Omega_{T}^{*}$ because of Condition 1.
Now we fix $\bm{\eta}\in\mathrm{relint}(\Delta^{K})$ and note that
$R_{L_{\Omega_{T}}}(\bm{\theta},\bm{\eta})=\Omega_{T}^{*}(\bm{\theta})+\mathbb{E}_{y\sim\bm{\eta}}[\Omega_{T}(\bm{\rho}(y))]-\langle\bm{\theta},\mathbb{E}_{y\sim\bm{\eta}}[\bm{\rho}(y)]\rangle$
is differentiable and convex in $\bm{\theta}$. If $\bm{\theta}^{*}\in\mathbb{R}^{\varrho}$ is the minimizer of $R_{L_{\Omega_{T}}}(\cdot,\bm{\eta})$, its optimality condition indicates
$\nabla_{\bm{\theta}}R_{L_{\Omega_{T}}}(\bm{\theta}^{*},\bm{\eta})=\nabla\Omega_{T}^{*}(\bm{\theta}^{*})-\mathbb{E}_{y\sim\bm{\eta}}[\bm{\rho}(y)]=0.$
When $\bm{\eta}\in\mathrm{relint}(\Delta^{K})$, we have $\mathbb{E}_{y\sim\bm{\eta}}[\bm{\rho}(y)]\in\mathrm{relint}(\mathrm{conv}(\bm{\rho}(\mathcal{Y})))$, and thus $\bm{\theta}^{*}$ satisfying the above optimality condition does exist. Finally, we conclude the proof by combining Lemmas 8 and 12.
Appendix B Detailed Proofs in Section 4
B.1 Proof of Corollary 18
Proof.  Denote by $V_{\bm{\theta}}(\bm{\pi})\coloneqq\sum_{i=1}^{K}e^{\theta_{i}+1-\pi_{i}}$. Since $\ln(\cdot)$ is strictly increasing on $\mathbb{R}_{>0}$, $V_{\bm{\theta}}(\bm{\pi})$ shares the same minimizer as $\ln\bigl{(}\sum_{i=1}^{K}e^{\theta_{i}+1-\pi_{i}}\bigr{)}$. The Lagrangian $\mathcal{F}$ of the minimization problem $V_{\bm{\theta}}$ for $\bm{\pi}\in\Delta^{K}$ is written as follows:
$\mathcal{F}(\bm{\pi},\bm{\alpha},\beta)=\sum_{i=1}^{K}e^{\theta_{i}+1-\pi_{i}}-\bm{\alpha}^{\top}\bm{\pi}+\beta\left(\sum_{i=1}^{K}\pi_{i}-1\right),$
where $\alpha_{1},\dots,\alpha_{K}\geq 0$ and $\beta\geq 0$ are the Lagrangian multipliers. Since $V_{\bm{\theta}}(\bm{\pi})$ is convex and differentiable, and the feasible region $\Delta^{K}$ is convex, the KKT conditions are necessary and sufficient for the optimality of $\bm{\pi}^{*}\in\Pi(\bm{\theta})$, which requires that there exists $(\bm{\alpha}^{*},\beta^{*})$ satisfying
$-e^{\theta_{y}+1-\pi_{y}^{*}}-\alpha_{i}^{*}+\beta^{*}=0,$ for any $y=1,\dots,K$, (20)
$\bm{1}^{\top}\bm{\pi}^{*}=1,\quad\bm{\pi}^{*}\geq 0,\quad\bm{\alpha}^{*}\geq 0,$ (21)
$\alpha^{*}_{y}\pi_{y}^{*}=0,$ for any $y=1,\dots,K$. (22)
The conditions (21) and (22) indicates that $\pi_{y}^{*}>0\implies\alpha_{y}^{*}=0$. Then according to (20), we have
$\pi_{y}^{*}>0\implies\pi_{y}^{*}=\theta_{y}-(\ln\beta^{*}-1).$ (23)
Meanwhile, for $\pi_{y}^{*}=0$, (22) indicates that $\alpha_{y}^{*}\geq 0$, which further indicates
$\pi_{y}^{*}=0\implies\theta_{y}=\ln(\beta^{*}-\alpha_{y}^{*})-1\leq\ln\beta^{*}-1.$ (24)
Then (23) and (24) can be rewritten as follows:
$\pi_{y}^{*}=\max\{\theta_{y}-(\ln\beta^{*}-1),0\}.$ (25)
According to (21) and $\ln\beta^{*}-1\in\mathbb{R}$, the KKT conditions can be further simplified into the existence of $\tau^{*}\in\mathbb{R}$ such that the following conditions simultaneously hold:
$\begin{cases}\pi_{y}^{*}=\max\{\theta_{y}-\tau^{*},0\},\\ \sum_{y=1}^{K}\max\{\theta_{y}-\tau^{*},0\}=1.\end{cases}$
Denote by $f(\tau)\coloneqq\sum_{y=1}^{K}\max\{\theta_{y}-\tau,0\}$. Then $f$ is continuous on $\mathbb{R}$, and moreover, it is strictly decreasing on $(-\infty,\theta_{[1]}]$ and equals 00 for $\tau\geq\theta_{[1]}$ because of the following expression:
$f(\tau)=\begin{cases}\sum_{y=1}^{K}(\theta_{y}-\tau)&\text{if $\tau\leq\theta_{[K]}$,}\\ \sum_{i=1}^{k-1}(\theta_{[i]}-\tau)&\text{if $\tau\in[\theta_{[k]},\theta_{[k-1]}]$ for $k=2,\dots,K$,}\\ 0&\text{if $\tau>\theta_{[1]}$}.\end{cases}$
Then $n$ defined in line 2 of Algorithm 1 exists since $f(\theta_{[1]})=0<1$. We also have that $f(\theta_{[n]})=\sum_{i=1}^{n}(\theta_{[i]}-\theta_{[n]})<1$. When $n<K$, we have that $f(\theta_{[n+1]})=\sum_{i=1}^{n+1}(\theta_{[i]}-\theta_{[n+1]})=\sum_{i=1}^{n}(\theta_{[i]}-\theta_{[n+1]})\geq 1$ by definition. Then we see
$\frac{\sum_{i=1}^{n}\theta_{[i]}-1}{n}\in[\theta_{[n+1]},\theta_{[n]}]$
and
$f\left(\frac{\sum_{i=1}^{n}\theta_{[i]}-1}{n}\right)$ $=\sum_{i=1}^{n}\left(\theta_{[i]}-\frac{\sum_{i=1}^{n}\theta_{[i]}-1}{n}\right)$
$=\sum_{i=1}^{n}\theta_{[i]}-\sum_{i=1}^{n}\theta_{[i]}+1$
$=1.$
When $n=K$, we have that $\sum_{i=1}^{K}\theta_{[i]}-K\theta_{[k]}<1$, that is,
$\frac{\sum_{i=1}^{K}\theta_{[i]}-1}{K}<\theta_{[K]}.$
Then
$f\left(\frac{\sum_{i=1}^{K}\theta_{[i]}-1}{K}\right)$ $=\sum_{i=1}^{K}\left(\theta_{[i]}-\frac{\sum_{i=1}^{K}\theta_{[i]}-1}{K}\right)$
$=\sum_{i=1}^{K}\theta_{[i]}-\sum_{i=1}^{K}\theta_{[i]}+1$
$=1.$
Combining these two cases, we have
$\tau^{*}=\frac{\sum_{i=1}^{n}\theta_{[i]}-1}{n}<\theta_{[1]},$
which is what line 3 of Algorithm 1 returns. Since $f(\tau)$ is strictly decreasing on $(-\infty,\theta_{[1]}]$, $\tau^{*}$ uniquely satisfies $f(\tau^{*})=1$.
Since $\sum_{i=1}^{k}\theta_{[i]}$ can be calculated cumulatively in $\mathcal{O}(K)$ time and the quick sort runs in $\mathcal{O}(K\log K)$ time, Algorithm 1 runs in $\mathcal{O}(K\log K)$ time in total. Furthermore, since $\tau^{*}$ is unique and $\pi_{y}^{*}=\max\{\theta_{y}-\tau^{*},0\}$, we ensure that $\Pi(\bm{\theta})$ is a singleton, which concludes the proof.
B.2 Proof of Proposition 19
Proof.  By Definition 10, the statement is equivalent to
$\mathop{\mathrm{missing}}{argmax}_{t\in\mathcal{Y}}\theta_{t}=\mathop{\mathrm{missing}}{argmax}_{t\in\mathcal{Y}}{{\pi}}_{\log}(\bm{\theta})_{t}.$
According to Corollary 18 and Algorithm 1, there exists $\tau\in\mathbb{R}$ such that $\pi_{\log}(\theta)_{i}=\max\{\theta_{i}-\tau,0\}$. We also have that $\max_{t}\theta_{t}>\tau$ by contradiction: if $\max_{t}\theta_{t}\leq\tau$, $\bm{\pi}_{\log}(\bm{\theta})=\bm{0}$ holds, which contradicts $\bm{\pi}_{\log}(\bm{\theta})\in\Delta^{K}$.
Denote the set $\mathrm{argmax}_{t\in\mathcal{Y}}\theta_{t}$ by $\mathcal{I}$. For any $i\in\mathcal{I}$, we have $\pi_{\log}(\bm{\theta})_{i}=\max\{\theta_{i}-\tau,0\}=\max_{t}\theta_{t}-\tau>0$. For any $j\not\in\mathcal{I}$,
  • If $\theta_{j}>\tau$: $\pi_{\log}(\bm{\theta})_{j}=\max\{\theta_{j}-\tau,0\}=\theta_{j}-\tau<\max_{t}\theta_{t}-\tau$,
  • If $\theta_{j}\leq\tau$: $\pi_{\log}(\bm{\theta})_{j}=\max\{\theta_{j}-\tau,0\}=0<\max_{t}\theta_{t}-\tau$.
These imply that for any $i\in\mathcal{I}$, $\pi_{\log}(\bm{\theta})_{i}\geq\pi_{\log}(\bm{\theta})_{j}$ and the equality holds if and only if $j\in\mathcal{I}$, which concludes the proof.